### v1

# Can Large Language Models Detect Misinformation in Scientific News Reporting?

Abstract.
Scientific facts are often spun in the popular press with the intent to influence public opinion and action, as was evidenced during the COVID-19 pandemic. Automatic detection of misinformation in the scientific domain is challenging because of the distinct styles of writing in these two media types and is still in its nascence. Most research on the validity of scientific reporting treats this problem as a claim verification challenge. In doing so, significant expert human effort is required to generate appropriate claims. Our solution bypasses this step and addresses a more real-world scenario where such explicit, labeled claims may not be available. The central research question of this paper is whether it is possible to use large language models (LLMs) to detect misinformation in scientific reporting. To this end, we first present a new labeled dataset SciNews, containing 2.4k scientific news stories drawn from trusted and untrustworthy sources, paired with related abstracts from the CORD-19 database, a comprehensive collection of COVID-19 and historical coronavirus research. Our dataset includes both human-written and LLM-generated news articles, making it more comprehensive in terms of capturing the growing trend of using LLMs to generate popular press articles. Then, we identify dimensions of scientific validity in science news articles and explore how this can be integrated into the automated detection of scientific misinformation. We propose several baseline architectures using LLMs to automatically detect false representations of scientific findings in the popular press. Our first architecture, SERIf, uses three modules: Summarization, Evidence Retrieval, and Inference to classify the news article as fake or true; the second architecture, SIf, bypasses the explicit evidence retrieval module while keeping the other two, and the third, direct-to-inference architecture, D2I, where both the summarization and the explicit evidence retrieval module are removed. For each of these architectures, we use several prompt engineering strategies including: zero-shot, few-shot, and chain-of-thought prompting. We also test these architectures and prompting strategies on GPT-3.5, GPT-4, and Llama2-7B, Llama2-13B. We note that the SIf architecture does best while the naive D2I does worst in terms of accuracy and F-1 scores.
Misinformation in Scientific Reporting, Large Language Models, AI-generated Misinformation, Explainability
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processingccs: Information systems Data mining
1. Introduction
Scientific information is communicated to the non-expert audience via popular press (news articles) and online platforms like blogs, social media posts, etc. Studies have shown that news with scientific-sounding content is trusted more than other types (Sharon and Baram-Tsabari, 2020). Therefore, any misinformation in the scientific domain can cause significant public risk as was evidenced during the recent COVID-19 pandemic (Mheidly and Fares, 2020). Other examples include the emergence of vaccination hesitancy (Razai et al., 2021; Baines et al., 2021), eroded trust in health institutions (Rodriguez-Morales and Franco, 2021), and the amplification of public fear and anxiety (Nelson et al., 2020; Verma et al., 2022). Therefore, it is imperative to identify misinformation in scientific news reporting.
Several websites are maintained by science reporters (Health News Review [^1] https://www.healthnewsreview.org/) and scientists (Science Feedback  [^2] https://science.feedback.org/) to track scientific misinformation in the media. Although such manual debunking is important, the sheer volume of scientific news can make this task unscalable.
Natural language processing (NLP) based approaches have consequently started to emerge to deal with this problem. These methods typically involve language analysis, like detecting exaggeration (Wright and Augenstein, 2021a) and certainty (Pei and Jurgens, 2021) and fact-checking (Guo et al., 2022a) and claim verification (Pradeep et al., 2020). Several claim verification datasets have also been developed for this problem (Thorne et al., 2018; Wadden et al., 2022) and a method for modeling information change from scientific article to scientific reporting has also been proposed (Wright et al., 2022a).
While these works have laid the foundation to address this problem, the area is still in its nascence. Several challenges remain unaddressed: 1) there are no existing taxonomies to define dimensions of the scientific validity of scientific news articles that can be used in automated methods to detect misinformation in scientific news; 2) all existing datasets for scientific fact-checking relies on explicit claim generation from the news articles before it can be compared to the scientific articles for misinformation detection. This can be a cumbersome process in real-life scenarios where it would be potentially necessary for expert human involvement to first generate claims from the scientific article and 3) there is no generalized architecture that can detect scientific misinformation without an explicit claim generation step.
In response to the aforementioned limitations, and because of the quantum leap in performance improvement offered by large language models (LLMs) in downstream NLP tasks, we formulate the following research questions:
  • RQ1: Can LLMs be used to define a general architecture to detect misinformation in scientific news reporting in simulated real-life scenarios without the need for explicit claims?
  • RQ2: Is it feasible to define dimensions along which the scientific validity of the news article can be measured and aid in the creation of effective prompts for these architectures?
  • RQ3: Do these architectures possess the capability to provide explanations for their decision-making processes?
To answer the above questions, we first create a novel dataset, SciNews [^3] We will make our dataset fully accessible to the research community, comprised of scientific news and related scientific articles. Given the rising trend in LLM-generated content in both legitimate reporting and misinformation, this dataset contains an equal number of LLM-generated and human-written articles. Our dataset construction method does away with extensive human annotation which can prove to be very expensive. The dataset construction pipeline is flexible enough to allow continuous updates with emerging news articles and scientific articles.
We then propose three architectures using LLMs to automatically detect false representations of scientific findings in the popular press without explicit claim generation. The first architecture, SERIf, uses three modules: Summarization, Evidence Retrieval, and Inference to classify the news article as fake or true; the second architecture, SIf, bypasses the explicit evidence retrieval module while keeping the other two, and the third, direct-to-inference architecture, D2I, adopts a direct-to-inference approach, dispensing with both summarization and explicit evidence retrieval. For each of these architectures, we employ several prompt engineering strategies including zero-shot, few-shot, and chain-of-thought prompting. We test these architectures using several state-of-the-art LLMs, including GPT (versions GPT-3.5 and GPT-4) and Llama2 (models with 7B and 13B).
Our extensive experiments conducted on the SciNews dataset have yielded significant findings: 1) we discovered that false scientific content generated by large language models (LLMs) is more challenging to detect compared to misinformation written by humans. 2) the SIf architecture arises capable of matching, and in some cases, surpassing the performance of the SERIf architecture. This suggests the feasibility of using LLMs to develop a more generalized architecture for detecting scientific misinformation more appropriate for real-life situations. 3) We make the first attempt to define dimensions of scientific validity and incorporate them into the Chain-of-Thought prompt. This novel addition has been instrumental in enhancing the LLMs’ ability to make more accurate predictions. 4) LLMs can generate plausible explanations for their decision-making process, which can help users understand the model’s judgment of news veracity.
2. Related Work
As mentioned earlier, scientific misinformation detection is still in its nascence and while related to misinformation detection in general, it is a harder problem since the language characteristics of the scientific communication is different from the formal format of scientific publications. The problem of scientific misinformation is related to two other concepts in NLP, including: 1) fact-checking (claim verification) 2) scientific language analysis.
However, none of these approaches can singly capture the complexity of scientific misinformation and so far, there has not been any attempt to systematically capture the ways in which scientific misinformation can occur and then to use that to detect scientific misinformation. In this work, we first define some dimensions of scientific validity and then harness the power of LLMs to design general architectures to analyze scientific news for misinformation.
2.1. Fact-Checking
Automatic fact-checking, which assesses the truthfulness of claims made in the text (Vlachos and Riedel, 2014; Guo et al., 2022b), has been extensively studied across various domains, including common knowledge verification (Thorne et al., 2018), political topics (Wang, 2017), COVID-19 (Saakyan et al., 2021), E-commercial (Zhang et al., 2020), biology (Chen et al., 2022). When thinking about scientific misrepresentation in popular media, it is natural to think of the veracity of scientific findings. It is therefore possible to cast the scientific misinformation as a fact-checking problem or claim verification problem. Several researchers have taken this approach to defect misinformation in the scientific domain (Vladika and Matthes, 2023; Wadden et al., 2020; Diggelmann et al., 2020; Wadden and Lo, 2021; Wadden et al., 2022; Wright et al., 2022b; Wang et al., 2023). These works typically construct claims from the existing scientific literature by manually reformulated scientific findings and then the constructed claims are verified by utilizing pre-existing knowledge resources. However, most of these works rely on human resources to identify and extract appropriate claims for verification. Furthermore, these artificially constructed claims may not accurately represent the complexity and nuance of claims encountered in real-world scenarios.
2.2. Scientific Language Analysis
Scholarly document processing has garnered considerable attention in recent years, reflecting a growing interest in the analysis and interpretation of scientific literature (Chandrasekaran et al., 2020). Of particular relevance to our research are tasks that track the change of scientific information from published literature to social press. This includes investigating writing strategies employed in science communication  (Tan and Lee, 2014; August et al., 2020), detecting changes in certainty (Pei and Jurgens, 2021) and exaggeration detection (Wright and Augenstein, 2021b). Furthermore, the automatic detection of semantic similarities between scientific texts and their paraphrases represents an alternative approach for analyzing scientific content (Lo et al., 2019; Piskorski et al., 2023). However, rather than use the typical metrics for measuring semantic similarity, we propose to use the inherent knowledge in pre-trained LLMs for this task.
2.3. Large Language Model in Misinformation
LLMs have consistently demonstrated the ability to generate text on par with human authors  (Zhang et al., 2023a; Yang et al., 2023). This has led to their widespread use by professionals in generating legitimate real news stories. Unfortunately, they have also been used to generate misinformation (Zhou et al., 2023; Chen and Shu, 2023; Buchanan et al., 2021) and often at a much larger scale than is humanly possible. While falsehoods crafted by LLMs prove challenging for humans to detect, compared to human-generated ones (Chen and Shu, 2023), several studies have illustrated the feasibility of identifying LLM-generated text using  (Tang et al., 2023; Buchanan et al., 2021). Motivated by these studies, we include a balanced set of LLM-generated scientific articles, both fake and true, in the SciNews dataset.
3. SciNews Dataset Construction
Figure 1. The dataset construction process: ① utilizing publicly available datasets as well as web resources to collect human-written scientific news related to COVID-19 (Subsection 3.1), ② selecting abstracts from CORD-19 as resources to guide LLMs to generate articles using jailbreak prompt (Subsection 3.2), ③ the dataset is augmented with evidence corpus drawn from CORD-19 (Subsection 3.3).
The first part of this work is to create a dataset, SciNews, of scientific news articles and associated technical scientific publications. The SciNews Dataset contains 2,400 news articles which are labeled: Reliable or Unreliable depending on whether the article represents scientific fact truthfully or not, respectively. The dataset contains 1,200 articles in each category to keep it balanced. Each article is systematically paired with up to three pertinent scientific abstracts from the CORD-19 repository. The CORD-19 is a comprehensive resource of over 1 million scholarly articles, including over 300,000 with full text, about COVID-19, SARS-CoV-2, and related coronaviruses (Wang et al., 2020).
Given the growing trend in using Large Language Models (LLMs) to generate legitimate as well as fake news stories (Zhou et al., 2023; Chen and Shu, 2023), we include LLM-generated news articles in the dataset. The SciNews dataset contains an equal number (1,200) of human-written and LLM-generated news articles in each category (reliable and unreliable). The statistics of the SciNews dataset are presented in Table 1. We show the construction of SciNews in Figure 1 and describe the construction of the dataset below.
Table 1. Distribution of article labels in SciNews Dataset.
Human-Written LLM-Generated Total
Reliable 600 600 1200
Unreliable 600 600 1200
Total 1200 1200 2400
3.1. Human-Written News Articles
To gather human-written news articles, we searched for content containing scientific information in existing misinformation datasets and known websites.
3.1.1. Leveraging Publicly Available Dataset
We leveraged MMCoVaR (Chen et al., 2021), COVID19-FNIR (Saenz et al., 2021), and COVID-Rumor (Cheng et al., 2021), which are labeled datasets containing human written news articles on COVID-19 from January 2020 through May 2021.
Our search within these datasets commenced with a predefined set of scientific keywords: {scientist, investigating, study finds, experts say, experts recommend}. Using these, we filtered the data to yield 1,190 candidate news pieces. Next, we manually reviewed each candidate to sift out articles without scientific content or irrelevant to COVID-19. The reason we eliminated articles that are irrelevant to COVID-19 was because we will be including scientific data from the CORD-19 as evidence. This process resulted in 223 news articles: 130 reliable and 93 unreliable.
3.1.2. Web-Based Collection
In order to expand the dataset to cover the latest discussions on COVID-19, we crawled both credible and dubious websites for data. To collect unreliable data, we referred to Wiki Fake News Website List [^4] https://en.wikipedia.org/wiki/List_of_fake_news_websites, crawled the listed sites for articles, and manually verified the content. We eliminated articles exhibiting blatant discrimination or prejudice and conspicuous propaganda devoid of substantial scientific dialogue. This process yielded 507 unreliable articles that contained discussions pertaining to COVID-19 and were grounded in a scientific context.
For reliable data, we restricted the range of sources for the news articles to a set of educational press sites and other well-regarded news websites. Appendix A.1 lists all the educational press sites. The full list of known trustworthy websites we consulted is included in Appendix A.2
The target news articles were collected by web crawling, anchoring our search with our set of scientific keywords augmented by two topical ones: COVID-19 and Vaccine. Each article was reviewed by the same set of annotators, ensuring a direct correlation with the referenced research papers. The content is scraped from the web to extract body text, title, and other data needed for data construction. We gathered 470 reliable articles from varied reputable sources in this way. This process gave us a combined total of 1,200 human-written news articles (600 reliable, 600 unreliable) spanning from January 2020 to October 2023.
3.2. LLM-Generated News Articles
Generating true articles using LLMs is fairly straightforward. However, since most LLMs come with guardrails to protect them from misuse, we used a jailbreak strategy to generate scientific misinformation. The step-by-step process to generate news articles using LLMs is described below.
3.2.1. Selecting Scientific Abstracts
First, we curate a collection of abstracts from the CORD-19 database. These abstracts served as the foundational resource for the subsequent generation process. The CORD-19 database contains more than 1,000,000 articles in the medical field. Typically, most widely distributed science articles tend to get more media attention. Hence, we focused on the most frequently cited papers from CORD-19. The CORD-19 organizes its articles using seven principal elements including {title, abstract, doi, PubMed ID, PMCID, JSON file ID, and XML ID}. We start our curation with papers that have all 7 elements. We then confine our focus to post-January 2020, subsequently filtering out off-topic data using the keyword set: {COVID-19, Corvarius, and Vaccine}. To further ensure that generated articles in our dataset are of high quality, we narrowed our selection to articles published in well-regarded journals spanning a spectrum from basic science (examples include ‘Cell’ and ‘Nature’) to medicine (such as the ’British Medical Journal’). The comprehensive list of these journals is detailed in Appendix A.3. From this refined pool, we handpicked the abstracts of over 2,000 highly cited articles, thereby forming a rich and diverse foundation for our dataset.
3.2.2. Jailbreak Prompts
We utilized the curated collection of abstracts as guiding resources for LLMs to generate both reliable and unreliable scientific articles. To mitigate the risk of generating harmful content, LLMs are subjected to an alignment process, complemented by the setup of predefined prompts serving as security measures (Achiam et al., 2023; Qi et al., 2023; TermsOfUseBing, 2023). Therefore, requests for LLMs to generate fake messages are usually denied. In light of these constraints and inspired by the Jailbreak Attacks (Mowshowitz, [n. d.]; Liu et al., 2023b, a; Shen et al., 2023), we designed a ‘Jailbreak’ prompt to enable the LLMs to generate fake news-oriented scientific articles that were both informative and contextually aligned with the provided abstracts. The designed jailbreak prompt is illustrated in Figure 2.
Figure 2. Schematic of the designed jailbreak prompt.
Given selected scientific abstract resources, we simulated a scenario where we acted as the instructor of a science class, prompting the large language model (LLM) to create two types of articles: a ‘True Article’ and a ‘Convincing False Article’. The objective was to use these articles as teaching tools to help students discern between authentic scientific knowledge and fabricated information. By including both true and false LLM-created content we can ensure that the systems trained on our dataset to detect false content are not simply detecting LLM-generated content, but would be adept at distinguishing between true and false content. The generated articles are generally in the style of a news article, with many including an explicit title, to mimic human-generated scientific news. In consideration of the cost implications associated with different LLMs [^5] https://openai.com/pricing666llama.meta.com, we primarily utilized Llama2-7B for generating the bulk of data samples, supplemented by a smaller set from GPT-3.5. After filtering (see subsection 3.4), this approach led to the creation of 1,000 data samples from Llama2-7b and 200 from GPT-3.5, reflecting a balance between quality and resource optimization.
3.3. Evidence Corpus Creation
To augment the constructed dataset, we matched as many as three scientific abstracts per news article as evidence resources. For both Human-Written and LLM-generated News Articles, we employed Vespa [^7] https://cord19.vespa.ai/ to identify relevant abstracts from CORD-19 based on BM25 scoring for each article. While most articles were matched with three corresponding abstracts, a few could only be paired with two or even just one. This led to the creation of a fixed evidence corpus comprising 7,087 pieces of paragraph-level evidence. While this evidence corpus remains static at this juncture, its design allows for future expansion.
3.4. Quality Control
The quality control team consists of 4 graduate students and 5 senior researchers with a background in NLP. For the Human-Written News Article subset and LLM-Generated subset, the team used different strategies to examine the data quality.
For the collection of human-written news articles from various sources, we referred to the guidelines outlined in  (Pustejovsky and Stubbs, 2012). Based on its principles and our specific needs, we developed an instruction guide, which can be found in Appendix A.5. To ensure uniformity and understanding of the task, all team members thoroughly reviewed this guide. An additional layer of quality assurance involved cross-checking the collected data among team members. This step was implemented to mitigate any potential biases and to guarantee that the data aligned with our collection criteria.
[…]
4. Dimensions of Scientific Validity and Proposed Architectures
Figure 3. Proposed Architectures. SERIf includes all three modules: Summarization, Sentence-level Evidence Retrieval, and Inference Module. SIf bypasses the evidence retrieval module while keeping the other two. D2I removes both the summarization and the explicit evidence retrieval module.
In order to develop automated methods to detect scientific misinformation in real-world situations, we first develop dimensions of scientific validity. These will later be used in the chain-of-thought prompts for the LLMs.
4.1. Dimensions of Scientific Validity
We define the following dimensions of scientific validity in the reporting. We note that this is not an exhaustive list of ways in which scientific validity may be compromised, however, to the best of our knowledge, this is the first attempt to systematically define directions of scientific validity in science news reporting for the design of automated science misinformation detection.
  1. (1)
    Alignment: The news paragraph may show different levels of alignment with the evidence sentences. Alignment in this case is defined as news and evidence representing the same meaning about one scientific content.
  2. (2)
    Causation confusion: The news article may confuse correlations presented in the scientific literature as causation. This could be one dimension in which the scientific validity is compromised.
  3. (3)
    Accuracy: This refers to how accurately the news item describes the scientific findings quantitatively and qualitatively
  4. (4)
    Generalization: This refers to overgeneralization or oversimplification of the findings reported in the scientific literature.
  5. (5)
    Contextual Fidelity: Does the news article retain the broader context of the scientific finding?
4.2. Proposed Architectures
Conceptually, we may think of the process of automatically detecting scientific misreporting (or mis/disinformation in science news reporting) “in the wild” as comprising of three elements: (1) understanding the gist of the news article; (2) comparing it to relevant information from scientific articles and (3) inferring if the news is reliable or unreliable. To this end, we propose three architectures with varying degrees of granularity. These architectures use LLMs and several prompting strategies for the different modules. Note that we do not require a separate claim generation module in any of the architectures. These architectures are depicted in Fig. 3.
In order to account for potential differences in performance between different prompting strategies and LLMs, each of these architectures are tested against multiple prompting strategies and LLMs and the results are described in Section 5.
4.2.1. The SERIf Architecture
The first proposed architecture, Summarization Evidence Retrieval Inference (SERIf), treats each of the conceptual elements above, as a separate module. Hence, it contains three modules: 1) Summarization; 2) Evidence Retrieval and 3) Inference. The summarization module distills the key information from the news articles, thereby streamlining the analysis process. The Evidence Retrieval module is responsible for identifying and extracting sentences from the scientific articles in our dataset that may validate or contradict the statements in the news article. This process aids in gathering relevant contextual evidence for further analysis. The Inference module categorizes the news articles into two reliable or unreliable, based on the evidence from the scientific dataset.
4.2.2. Summarization module
Inspired by the recent success of text summarization (Zhang et al., 2023b) using LLMs, we take the ’Extractive - Abstractive’ two-step summarization strategy (Zhang et al., 2023c) to construct a summary of the news article. The extractive summarization process summarizes the article by identifying and concatenating the most salient sentences from it, ensuring that the extracted summaries are consistent with the original text. The resulting summaries serve as a foundation for the abstractive summarization which uses a generative approach to create a more concise and cohesive summary. By synthesizing the extractive and abstractive approaches, the module ensures a balance between accuracy (adherence to the original text) and brevity (conciseness and essence of the content), thereby providing an effective and reliable summarization for further analysis in the misinformation detection process.
Formally, for a document composed of $n$ sentences, the extractive summarization process creates an extractive summary, $S_{e}$, consisting of $m\ll n$ sentences. Then, the LLM, $M$, creates an abstractive summary using a query, $q$ and $S_{e}$ as input: $S_{a}=M(q,S_{e})$.
As a sanity check, to verify the quality of the summary, we randomly selected 200 samples for annotation by two graduate students with a background in NLP, and evaluated the summaries using four criteria: 1) quality of extractive summary (High vs Low); 2) quality of abstractive summary (High vs Low); 3) presence of hallucination in abstractive summary (Yes vs No); 4) comparison of Extractive and Abstractive Summaries. We used the Krippendorff’s alpha score to evaluate the agreement between the annotations (Hayes and Krippendorff, 2007). The alpha scores for the four aspects were 0.53, 0.82, 0.91, and 0.94, respectively. These values suggest that there is strong agreement that the abstractive summaries effectively encapsulate the core information of the original texts and maintain a high degree of consistency. Consequently, we have decided to use the abstractive summaries in subsequent steps of our analysis.
4.2.3. Sentence-level Evidence Retrieval
The process of key evidence selection in our The evidence retrieval module involves extracting key sentences from scientific articles that may support or refute the claims of the news article. This task bears a resemblance to paragraph retrieval but operates at a finer granularity. It essentially constitutes a semantic matching challenge, where each sentence within a paragraph undergoes a comparison against a specific statement query. The objective is to pinpoint the most relevant evidence interval within these sentences.
A critical step in refining this process was pre-defining our evidence corpus using CORD-19. This strategic choice significantly narrows down the search space to a manageable scope, allowing for efficient traversal through all relevant paragraphs to locate key evidence. We use an LLM to enhance the effectiveness and accuracy of our sentence selection process.
Given an abstractive summary, $S_{a}$, and candidate scientific abstract, $E$, we select sentences $e_{i}$ from $E$: $\{e_{i}\}=M(S_{a},E)$, where $\{e_{i}\}$ is a set that contains all relevant and important sentences selected by LLM based on semantics.
4.2.4. Inference Module
The Inference Module is dedicated to assessing the the veracity of the summarized news paragraph using the set of retrieved evidence sentences using the abstractive summary, $S_{a}$, and the selected evidence sentence set $\{e_{i}\}$. Thus, the inference module produces a binary output (reliable or unreliable) for each $<S_{a},\{e_{i}\}>$ pair.
Table 2. Performance results of our proposed SERIf architecture using different LLMs and prompt strategies.
Models Prompt Strategy Human-Written LLM-Generated Overall
Accuracy Precision Recall F1 Accuracy Precision Recall F1 Accuracy Precision Recall F1
GPT-3.5 Zero-Shot 74.25 74.95 72.83 73.87 66.75 60.88 93.66 73.80 70.50 67.92 83.23 73.84
Few-Shot 70.00 71.64 66.60 68.95 68.14 62.12 92.82 74.43 69.07 66.88 79.71 70.49
CoT 76.67 76.20 77.67 76.92 66.75 60.66 95.33 74.14 71.71 68.43 86.50 75.53
GPT-4 Zero-Shot 77.33 76.48 79.00 77.72 70.25 62.91 98.67 76.83 73.79 69.70 88.34 77.26
Few-Shot 75.08 74.60 76.00 75.30 70.17 62.85 98.32 76.68 72.63 68.73 87.16 75.99
CoT 79.58 76.90 85.00 80.72 67.25 60.50 99.33 75.20 73.42 68.70 92.17 77.96
LLAMA2-7B Zero-Shot 56.00 53.24 98.33 69.10 51.17 50.60 96.80 66.50 53.89 51.92 97.57 67.80
Few-Shot 54.75 58.20 93.30 71.70 52.00 51.00 97.30 67.00 52.38 54.60 95.30 69.35
CoT 56.83 59.20 97.80 73.70 51.58 50.80 96.80 66.70 54.21 55.00 97.30 70.20
LLAMA2-13B Zero-Shot 57.33 59.50 98.50 74.20 53.67 57.10 97.00 71.90 55.50 58.30 97.75 73.05
Few-Shot 56.91 59.50 95.80 73.40 52.33 51.20 97.50 67.20 54.62 55.35 96.65 70.30
CoT 58.00 59.90 99.00 74.60 55.08 52.70 98.30 68.60 56.54 56.30 98.65 71.60
4.2.5. The SIf Architecture
In this architecture, we remove the evidence retrieval module from the previously described SERIf architecture. In the SIf architecture, the Summarization module works exactly as described in Section 4.2.1. The LLM in the Inference module is now directly prompted to decide whether the given news summary is trustworthy or not and to provide justifications based on the paired scientific abstracts from the evidence corpus in the SciNews dataset.
4.2.6. The Direct to Inference (D2I) Architecture
In the third architecture, there is no summarization module or explicit evidence retrieval module. Instead, the LLM is directly fed the scientific news article, and the corresponding scientific abstracts and prompted to determine whether the news item is trustworthy with justifications.
When viewed from the perspective of identifying scientific misinformation ”in the wild”, the D2I is the architecture that does little in the way of processing and the SERIf architecture involves the most processing. In other words, the SERIf requires engineering each aspect of the elements of scientific misinformation separately, the D2I architecture requires very little engineering and the SIf falls between these two. However, as noted earlier, none of these architectures expect an explicit set of claims to be generated from the news article for misinformation detection.
4.2.7. Prompt Strategies
A key factor in the performance of an LLM-based task is prompt engineering. Several kinds of prompting strategies have been used in various applications with varying degrees of success in specific tasks.
  • Zero-shot prompting: LLMs are presented with a task without any prior specific training or examples related to that task (Brown et al., 2020). We test the performance of LLM on scientific misinformation detection by zero-shot prompts directly using the pre-existing knowledge of the LLMs.
  • Few-shot prompting: In contrast to zero-shot prompting, where LLMs are presented with tasks without prior training or examples, few-shot prompting involves furnishing LLMs with a concise set of examples prior to task execution (Brown et al., 2020). This approach is designed to provide the model with essential context, thereby augmenting its capability for tasks like detecting scientific misinformation. In our methodology, we supply the LLMs with two carefully selected examples: one deemed ’reliable’ with accompanying reasoning, and another labeled ’unreliable’. This strategy is intended to better equip the LLMs to discern and categorize information accurately in the execution of the task.
  • Chain-of-thought prompting (CoT): CoT prompting involves structuring prompts to elicit a step-by-step reasoning process, effectively emulating the cognitive process humans employ in solving complex problems (Xu et al., 2022). In our approach, we used the dimensions of scientific validity defined in Section 4.1 to design multiple CoT prompts to guide the LLMs. This methodology not only aids the LLMs in systematically dissecting and assessing factual content but also aligns their reasoning process with structured, human-like analytical methods.
5. Results and Discussions
We used GPT-3.5, GPT-4, Llama2-7B, and Llama2-13B in our architectures with the temperature set to 00. This setting ensures that the LLMs generate responses with the highest predictability, thereby maintaining uniformity across the experiments. The performance of each of the proposed architectures, using each of the above LLMs is measured using accuracy, precision, recall, and F1-score. These results are displayed in Tables  2 and 3.
We note immediately that there is a significant difference between the performances of GPT and Llama2, with both GPT-3.5 and GPT-4 beating Llama2-7B/13B. All Llama2-7B/13B models achieved an accuracy score barely above random guessing. This may be due to the greater popularity of GPT, which may have made the crowdsourced, ongoing training of the GPT models stronger than those of Llama2.
5.1. Human-Written vs LLM-Generated Misinformation
Table 2 records the results of our experiments on the SERIf architecture and Table 3 shows the performance of the SIf and D2I architectures. From these tables, we note that there is a significant difference in misinformation detection accuracy and F1 scores for the human-written articles and LLM-generated articles. Across all architectures, it is consistently more challenging to identify LLM-generated scientific misinformation compared to human-authored misinformation. This is evidenced by high recall scores paired with low precision scores, indicating a tendency of the detectors to classify news as True. Such a trend highlights the difficulty in discerning false information within LLM-generated scientific news. This notable discrepancy emphasizes the complexity and challenge inherent in detecting misinformation in LLM-generated content. It aligns with similar findings in non-scientific misinformation domains (Chen and Shu, 2023). These results also raise concerns about the potential misuse of LLMs and underscore the importance of advancing our detection methodologies to keep pace with the evolving capabilities of LLMs, considering their implications for public safety.
Table 3. Performance results of our proposed SIf and D2I architectures using different LLMs and prompt strategies.
Models Arch Prompt Strategy Human-Written LLM-Generated Overall
Accuracy Precision Recall F1 Accuracy Precision Recall F1 Accuracy Precision Recall F1
GPT3.5 SIf Zero-Shot 82.50 80.00 86.50 83.50 62.00 57.00 99.00 72.00 72.25 68.50 89.13 77.75
Few-Shot 78.67 79.76 76.83 78.27 65.33 59.89 92.83 72.81 72.00 69.83 84.83 75.54
CoT 76.08 79.88 73.67 75.68 70.17 65.12 86.83 74.43 73.13 72.50 80.25 75.06
D2I Zero-Shot 66.60 66.55 64.33 65.42 63.42 57.85 98.33 72.98 65.01 62.20 81.33 69.20
Few-Shot 65.60 63.30 67.33 65.25 62.75 63.80 97.53 77.10 64.18 63.55 81.33 71.18
CoT 77.17 69.50 96.83 80.91 64.08 57.63 99.65 73.03 70.63 63.57 98.24 76.97
GPT-4 SIf Zero-Shot 80.00 80.00 79.00 80.00 71.00 64.00 98.00 77.00 75.50 72.00 88.50 78.50
Few-Shot 70.08 75.91 58.83 66.29 71.75 64.17 98.50 77.71 70.92 70.04 78.67 72.00
CoT 73.92 82.38 60.83 69.98 71.08 65.79 87.83 75.23 72.50 74.09 74.33 72.61
D2I Zero-Shot 68.08 66.80 72.00 69.30 65.00 59.00 98.50 73.80 66.50 62.90 85.25 73.80
Few-Shot 70.00 71.40 66.70 69.00 68.14 62.20 92.82 74.50 69.07 66.80 79.71 71.75
CoT 78.08 84.60 68.67 75.80 72.00 65.20 98.50 78.30 75.04 74.90 83.56 77.05
5.2. Comparison Across Architectures (RQ1)
The comparative analysis revealed a significant performance superiority of GPT over Llama2. Therefore, GPT was selected for further evaluation of the various architectures proposed in this work. According to the result presented in Table 3, it is evident that the SIf architecture yields the highest accuracy in processing human-authored articles, achieving an impressive $82.50\%$ in accuracy and $83.50\%$ in F1 score. The encouraging results, even in the absence of the ‘sentence-level evidence retrieval’ module, suggest the potential to develop more flexible and generalized scientific misinformation detection models. In stark contrast, the performances of the D2I models were notably subpar. This disparity indicates that merely supplying a Large Language Model (LLM) with a news article coupled with scientific articles does not suffice to guarantee effective performance. Furthermore, our results highlight the importance of the ‘summarization’ module. By distilling key statements from the news, this module minimizes the impact of extraneous information, thereby enhancing the LLM’s ability to generate more accurate predictions.
Table 4. Example of prompt and response for generating explanations.
Prompt: Decide if the evidence ’support’ or ’refute’ the point of view of the news paragraph. Provide your prediction and explain your reasoning in a few sentences.
Response: Refute. The evidence provided does not support the claim made in the news paragraph. While the evidence mentions neurological complaints reported by vaccinated individuals, they do not provide any evidence that suggests COVID-19 mRNA vaccines cause extensive brain damage. Additionally, the fact that different vaccines produced different profiles of complications does not necessarily indicate brain damage. The study’s authors calling for further investigations does not confirm the claim either.
5.3. Comparison Across Prompting Strategies (RQ2)
Observing the overall performance from Table 2 and 3, the results reveal a significant trend: the CoT prompting generally outperforms the zero-shot approach. Notably, for LLM-generated data, the CoT prompt markedly enhances detection capabilities. This suggests that our defined dimensions of scientific validity effectively aid LLMs in making more accurate predictions. However, an interesting observation in the SIf architecture for human-authored content is that, while the CoT prompt led to higher precision, it resulted in lower recall compared to the zero-shot approach. This indicates that our scientific validity dimensions sometimes prompt the LLM to apply more stringent criteria in its judgments.
On the other hand, the few-shot setting did not significantly improve performance, implying that despite providing two well-crafted examples (one positive and one negative), it is challenging for LLMs to extract substantial features from the provided cases. This not only highlights the complexity of scientific misinformation detection but also underscores the intricate nature of the potential scientific misinformation data involved.
5.4. Explainability Study (RQ3)
To assess the explainability of the proposed architectures, we prompt the LLMs to not only classify the news articles as reliable or unreliable, but also to explain the reasoning behind these classifications. An example of the prompt and the result are presented in Table 4. The response illustrates the LLM’s ability to the content of the news article against corresponding evidence and articulates the differences, leading to explanations that are comprehensible to users. This feature is crucial as it aids individuals in discerning the veracity of scientific content, enhancing their understanding of what constitutes true versus false science. We provide the complete example in the appendix C.
6. Conclusions
[…]
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.
Appendix A Materials to SciNews Dataset Construction
A.1. Educational Press Sites
We collected data from the following educational press sites:
YaleNews, Yale School of Medicine Latest News, Boston University – University News, Boston College BC News, University of Washington School of Medicine Newsroom, Regenstrief Institute News, University of South Carolina In the News, University of Utah Unews, Colorado State University Source, University of Kansas Medicine Center News, University of Michigan News, University of Nebraska Medicine News & Events, University of Maryland School of Medicine News, Stanford News, Stanford Medicine News Center, University of Mississippi Medical Center News Stories, Washington University School of Medicine in St. Louis News, Center for Education Policy Research at Harvard University News, Johns Hopkins Bloomberg School of Public Health Articles & News Releases, University of Missouri School of Medicine News, University of Hawaii News, The Ohio State University Wexner Medical Center Press Releases, Oregon State University Newsroom, University of Minnesota News and Events, Emory University Emory News Center, Tufts Now, University of Kentucky College of Medicine News, University of Calgary UCALGORY News, Texas AM Today, Duke Today, North Carolina State University NC State News, Vanderbilt University Research News, University of Toronto U of T News, McMaster University Daily News, University of Virginia UVAToday, University of New Hampshire Newsroom, Rutgers University – Rutgers Today, UT Southwestern Research Labs News, University of Houston UH Newsroom, University of Oxford News, Queen Mary University of London Queen Mary News, University of York News, The BMJ News, JAMA Health Medical News, Nature News, Allen Institute News, National Institutes of Health News and Events.
A.2. News Sources
We extracted data from the following news websites: CNBC, The Washington Post, The Atlantic, CNN, NPR, BBC, Forbes, USA Today, Bloomberg, Daily Mail, CBC, News Medical, ABC News, CBS News, The Economic Times, and OHSU News.
Although these sources are trustworthy on the whole, there can still be some biased content. Our team double-checked the content. Only after a rigorous verification process was an article deemed suitable for inclusion in our dataset.
A.3. Journal List
Academic articles in journals with good reputations are more likely to attract attention and be widely disseminated. Therefore, we select abstracts of high-quality articles from the CORD-19 database based on the following list to be used as resources for LLM-generated articles:
Nature, Science, British Medical Journal, Journal of Medical Virology, BMC Medicine, Blood, Nature Cell Biology.
Figure 4. An example of LLM-generated articles by using JailBreak prompt.
A.4. Review Guidance for Human-Written Articles
To ensure that our dataset covers only scientific-related content and does not include politics, economics, etc., we apply the following guidance to check each collected human-written article:
The article can include:
  • After reading through the title and body text, the main content is the discussion of scientific discoveries or scientific progress.
  • The title contains obvious scientific vocabulary: such as investigating, study finds, scientist, experts say, and ‘experts recommend’.
  • The title reads as a scientifically relevant conclusion or discussion:
  • The main body content is some news summary or news paraphrase.
The article cannot include:
  • Live News Style Title.
  • Explicit political information.
  • Contains other information such as finance and marketing in the title.
  • If First-person pronouns appear in the title, it should be noted that it is not a science-related discussion.
A.5. An Example of LLM-Generated Data Sample
Figure 4 showcases a Llama2-7B-generated LLM article example. The ’True Article’ mirrors the original scientific abstracts accurately, whereas the ’Convincing False’ modifies key details, including experimental effects and scientific conclusions. This indicates that even LLMs with smaller parameters, such as llama2-7B, are capable of producing scientific disinformation. Fabricating and exaggerating medical research findings could lead to public complacency in pandemic situations, resulting in greater harm. This example not only validates our jailbreak prompt approach but also underscores the public safety risks associated with LLMs.
Appendix B A Summary of Prompt Examples
In our experiments, we employed various prompt strategies: zero-shot, few-shot, and Chain-of-Thought (CoT). For the different architectures, the prompt remained consistent across all strategies.
Table 5. An example of the Zero-Shot prompt.
System Message: As a Fact Checker, your role involves analyzing a news paragraph and several evidence sentences provided by the user. The user will present a news paragraph. Following this, the user will present evidence sentences. Your task is to determine the factual accuracy of the news story based on these evidence sentences. To justify your conclusion, select and reference specific phrases or sentences from both the news story and the evidence provided.
User Message: I will give you one news paragraph and several relevant sentences. Please help me determine if these sentences support or refute the news point of view. Finally, please answer using one word ‘refute’ or ‘support’ and give reasons. Please provide the final output in JSON format containing the following two keys: prediction and reason.
Table 6. An example of the Few-Shot prompt.
System Message: As a Fact Checker, your role involves analyzing a news paragraph and several evidence sentences provided by the user. The user will present a news paragraph. Carefully read this paragraph to understand its central claim. Following this, the user will present evidence sentences. These sentences may either support or refute the news paragraph’s central claim. Your task is to determine the factual accuracy of the news story based on these evidence sentences. Are they supporting or contradicting the news? To justify your conclusion, select and reference specific phrases or sentences from both the news story and the evidence provided.
User Message: Task: Analyze the following news paragraph and several relevant sentences to determine their relationship.
Example 1: {One positive example with label and reason.}
Example 2: {One negative example with label and reason.}
Now analyze the following: {News Paragraph} and {Evidence Corpus}
Instructions: Decide if the relevant sentences ’support’ or ’refute’ the point of view of the news paragraph. Provide your answer in one word - either ‘support’ or ‘refute’. Then, explain your reasoning in a few sentences. Output: format your response as JSON with two keys: prediction and reason.
Table 7. An example of the Chain-of-Thought (CoT) prompt involves integrating our defined dimensions of scientific validity into the thought process, guiding the LLM toward making decisions.
System Message: You are a Fact Checker. The user will present a new paragraph. Following this, the user will present evidence paragraphs. These sentences may either support or refute the news paragraph’s central claim. Your task is to determine the factual accuracy of the news story based on these evidence paragraphs. Make a final prediction and provide a comprehensive explanation step by step based on the following:
Alignment Check: examine the evidence for alignment with the news paragraph
Causation confusion: evaluate if the news paragraph confuses correlation with causation
Accuracy: verify quantitative and qualitative accuracy in the news paragraph compared to evidence
Generalization: assess if the news paragraph overgeneralizes or oversimplifies findings from evidence sentences
Contextual Fidelity: consider the broader context surrounding the news and evidence.
User Message: I will give you one news paragraph and relevant evidence corpus. Please help me determine if these paragraphs support or refute the news point of view. Finally, please answer using one word ‘refute’ or ‘support’ and give reasons. Please provide the final output in JSON format containing the following two keys: prediction and reason.
Appendix C More Details of Explainability Study
The example in figure 5 is from SERIf architecture under the CoT prompt and GPT-4. It shows how the ’sentence-level evidence retrieval’ module effectively identifies relevant statements from the evidence corpus, detects contradictions between the original text and evidence, and makes accurate predictions during the ’Inference’ phase based on the predefined dimensions of scientific validity. Such effective explanations enhance understanding of the reasoning process.
Figure 5. An example of explainability study.